# SlumpGuard: An AI-Powered Real-Time System for Automated Concrete Slump Prediction via Video Analysis

### v2

@@ Abstract
Concrete workability is essential for construction quality, with the slump test being the most common on-site method for its assessment. However, traditional slump testing is manual, time-consuming, and prone to inconsistency, limiting its applicability for real-time monitoring. To address these challenges, we propose SlumpGuard, an AI-powered, video-based system that automatically analyzes concrete flow from the truck chute to assess workability in real time. Our system enables full-batch inspection without manual intervention, improving both the accuracy and efficiency of quality control. We present the system design, the construction of a dedicated dataset, and empirical results from real-world deployment, demonstrating the effectiveness of SlumpGuard as a practical solution for modern concrete quality assurance.
@@ 1 Introduction
Concrete stands as the backbone of modern construction infrastructure, with over 14 billion cubic meters produced annually worldwide, making it the most widely consumed material after water [1]. Among the critical properties governing concrete performance, workability—primarily assessed through slump testing—represents a fundamental parameter that directly influences mixing efficiency, transportation logistics, placement operations, and compaction effectiveness on construction sites [2, 3]. However, the complex and dynamic nature of construction environments presents unprecedented challenges for maintaining consistent concrete quality control. In particular, conventional manual slump testing, still the industry standard, often fails to meet the demanding requirements of large-scale, time-sensitive projects.
The construction industry therefore continues to struggle with real-time quality assurance. Traditional slump testing remains labor-intensive, operator-dependent, and unable to provide continuous monitoring during concrete placement. Manual testing introduces delays, increases costs, and is sensitive to environmental factors, while reliance on less experienced personnel further undermines consistency [4]. Although automated approaches such as video-based slump prediction, stereo vision, depth sensing, and IoT monitoring have been explored, most are validated only in controlled or small-scale settings and face challenges such as variable lighting, dust, and equipment interference in dynamic construction environments [5, 3, 6]. Moreover, these systems often depend on specialized hardware or require significant infrastructure modifications, limiting their practicality. Consequently, there is an urgent industry need for field-deployable, cost-effective solutions capable of providing robust, real-time monitoring of concrete workability without disrupting construction workflows, addressing environmental variability, and integrating seamlessly with existing operations [7, 8].
To address the limitations of conventional slump testing in real-site applications, we present SlumpGuard, a comprehensive AI-powered system specifically engineered for real-world deployment in construction environments. Unlike manual methods that are labor-intensive, inconsistent, and unsuitable for continuous monitoring, SlumpGuard operates directly on the flow of ready-mix concrete from mixer truck chutes to enable automated, full-batch inspection of every delivery without manual intervention. Our approach introduces several key innovations: (1) a robust three-stage pipeline combining oriented object detection, optical flow analysis, and video classification, (2) advanced data augmentation strategies specifically designed to handle the environmental variability encountered in construction sites, and (3) comprehensive validation using a dataset collected over eight months from actual construction conditions rather than laboratory simulations. This video-based system not only enhances the consistency and reliability of slump measurements but also streamlines quality assurance procedures across the entire construction workflow, ultimately contributing to safer, more efficient, and higher-quality construction outcomes.
@@
The remainder of this paper is organized as follows. Section 2 reviews existing studies on concrete workability assessment and automated monitoring approaches. Section 3 describes the dataset collection process and annotation methodology. Section 4 details the architecture of the proposed SlumpGuard system, including its three-stage pipeline and data augmentation strategies. Section 5 presents experimental results and performance evaluation under real-site conditions. Finally, Section 6 concludes the paper and outlines future research directions.
@@ 2.1 Optical Flow
Optical flow is a computer vision technique used to estimate the motion of objects or scenes across video frames by analyzing the movement of pixels. It represents motion as a vector field, where each vector indicates the displacement of a pixel between consecutive frames. There are two main approaches: dense optical flow, which computes motion for every pixel, and sparse optical flow, which estimates motion only at selected feature points. While dense optical flow [9] offers a detailed view of the entire scene’s motion, it is computationally intensive. Sparse optical flow is more efficient but may overlook fine-grained motion.
In this work, we adopt the Lucas-Kanade method [10], a widely used algorithm for sparse optical flow, suitable for real-time applications such as tracking concrete flow. It assumes that the intensity of a point remains constant over time:
@@
FlowNet [11] pioneered the use of convolutional neural networks for end-to-end optical flow prediction, learning motion patterns directly from synthetic datasets. Building on this idea, FlowNet2 [12], PWC-Net [13], and LiteFlowNet [14] introduced architectural improvements such as pyramid-based warping, cost volume reasoning, and hierarchical refinement. More recently, RAFT [15] achieved significant performance gains by leveraging dense, all-pairs correlations and iterative updates to produce highly accurate flow estimates.
Previous studies have applied optical flow combined with imaging techniques for slump prediction under controlled experimental settings [4, 5]. In contrast, our approach is specifically designed for robust, real-world deployment on construction sites, enabling continuous, automated slump quality monitoring under diverse and dynamic field conditions.
@@ 2.2 Object Detection
Object detection is a fundamental computer vision task widely applied in construction to automatically identify and localize key entities such as workers, equipment, and materials in images or videos. Recent studies have demonstrated its effectiveness in construction site monitoring [16, 17].
State-of-the-art detection frameworks such as Faster R-CNN [18], Mask R-CNN [19], SSD [20], and YOLO [21] offer a good balance between speed and accuracy, making them well-suited for real-time applications on construction sites.
However, in construction environments where objects like materials and equipment often appear at arbitrary angles, standard axis-aligned bounding boxes may be insufficient. In such cases, oriented object detection becomes essential. Early approaches like R2CNN [22] introduced rotation-invariant proposals, while more recent models such as Oriented R-CNN [23] and ReDet [24] employ rotation-aware architectures to improve detection accuracy. These methods are particularly effective for automated inspection tasks involving rotated or non-axis-aligned objects.
@@ 2.3 Video Classification
Video classification is a fundamental task in computer vision that involves assigning semantic labels to video clips by analyzing both spatial and temporal information inherent in the data. Recent advances in this domain have been largely driven by two primary classes of deep learning architectures: 3D convolutional neural networks (3D CNNs) and vision transformers [25] tailored for video.
3D CNNs like ResNet3D [26] extend 2D convolutions to the temporal dimension, effectively capturing spatiotemporal features and achieving strong results on benchmarks such as Kinetics-400 [27]. Transformer-based models, notably TimeSformer [28], use divided spatial and temporal self-attention to model long-range dependencies efficiently, surpassing many convolutional approaches. Further advancements include VideoSwin [29] and MViT [30], which incorporate hierarchical and multiscale attention mechanisms to improve video representation learning and classification performance.
These video classification models have also been increasingly applied in the construction industry, playing a pivotal role in streamlining quality supervision, activity recognition, and concrete workability assessment. Previous studies have also explored video-based approaches for material quality monitoring and evaluation [31, 3, 8], alongside behavior classification systems aimed at improving worker safety and productivity [32, 33].
@@ 3.1 Data Collection
In the experimental process, considering cost efficiency, we strategically utilized a single stereo camera for video acquisition. The stereo camera allowed simultaneous capturing of video footage from two distinct viewpoints in a single filming session, allowing richer data collection within a limited number of experiments. This approach maximized data efficiency while clearly capturing the concrete flow through the chute. Additionally, we performed traditional slump cone measurements at regular intervals, establishing precise ground-truth labels corresponding to each video frame. This example is shown in Figure 1.
@@ 3.2 Preprocessing and Annotation
Since we employed a stereo camera setup, the collected videos consisted of two separate streams. Therefore, as part of preprocessing, the original stereo data was divided into two individual video sequences, effectively doubling the amount of data for analysis. Considering that concrete discharges directly from the ready-mixed concrete chute, our preprocessing specifically targeted the chute region. Bounding box annotations were performed by categorizing into two classes, as shown in Figure 2, an “Unrotated Chute” representing the entire chute region with axis-aligned boxes, and a “Chute” represented by rotated bounding boxes.
@@ 4.1 Chute Detection
To automatically identify when concrete is falling from the chute, we first detect the chute region in the video using an object detection model. We employ YOLOv8 [34], a state-of-the-art object detection model, known for its fast inference performance and high accuracy, making it suitable for real-time object detection. We initialize the model with weights pretrained on the DOTA dataset [35], which is designed for detecting oriented objects in aerial imagery and is well-suited for our task of detecting rotated chutes. Using this model, we simultaneously predict both axis-aligned and rotated-bounding boxes for the chutes, enabling us to detect the chute area in pouring videos. The detected chute is represented by a bounding box with parameters $(x,y,w,h,\theta)$, indicating the center coordinates, width, height, and rotation angle of the box.
@@ 4.2 Detecting Concrete Placement Location and Timing
On construction sites, concrete is discharged from two separate chutes connected to individual concrete mixer trucks. In this study, we detect which chute the concrete is falling from and accurately identify the starting point of the flow. This step is essential to ensure that the subsequent video classification model receives only the segments where concrete is actually flowing, enabling more precise slump analysis. Since video segments without concrete flow are not relevant for analysis, filtering out such irrelevant data in advance is a critical preprocessing step.
To precisely determine when and from which chute the concrete starts to fall, we perform motion analysis based on the segmented chute videos $V^{L}$ and $V^{R}$ obtained in Step 1. For each video segment, we extract bounding boxes for the chutes from the object detector and compute optical flow at the center point of each bounding box. Specifically, we apply the Lucas-Kanade method, a sparse optical flow algorithm, to track the motion vector $(u,v)$ at the center $(x,y)$ across consecutive frames. When the tracked center point crosses the bottom boundary of the bounding box over time, we identify the moment of concrete drop as well as the specific chute through which the concrete is discharged.
@@ 4.2.2 Determining the Location and Timing
To make the temporal change explicit, we define the center point of the tracked chute in each frame as $(x_{t},y_{t})$ at time $t$, and as $(x_{t-1},y_{t-1})$ at the previous frame t1t-1. The optical flow vector $(u,v)$ represents the displacement from frame t1t-1 to $t$, so that $(x_{t},y_{t})=(x_{t-1}+u,y_{t-1}+v)$. The endpoints of the bottom edge of the rotated bounding box, $\mathbf{p_{1}}=(x_{1},y_{1})$ and $\mathbf{p_{2}}=(x_{2},y_{2})$, are computed.
@@ 4.3 Slump Prediction and Anomaly Detection
To enable efficient and real-time estimation of concrete slump, we leverage the drop timing and location identified in Section 4.2. Specifically, given the drop time $t$, we predict the slump using video data from the corresponding chute region. To focus on the continuous flow of concrete, we adopt a video classification model, ResNet 3D [26]. Let $V$ denote the chute video (i.e., $V^{R}$ or $V^{L}$), and $t$ the detected drop time.
@@ 4.3.2 Data Augmentation
Given that construction sites are exposed to a wide range of lighting and environmental conditions, data augmentation is essential to improve the model’s generalization performance. To simulate these real-world variations, we employ a set of augmentation techniques, including ColorJitter, gamma correction, horizontal flipping, and contrast adjustment. We also apply MixUp [36] and CutMix [37], which blend either the pixel content or the labels of two training examples to generate mixed samples. While MixUp interpolates both images and labels, CutMix replaces a region of an image with a patch from another image, adjusting the labels proportionally to the area. These strategies encourage the model to learn more robust and smoother decision boundaries, and to be less sensitive to spurious correlations between local image regions and class labels.
Together, these augmentations help the model handle diverse visual appearances–such as changes in brightness, shadows, occlusion, and orientation–that are commonly encountered in construction sites environment. They also promote better regularization by expanding the training distribution and reducing overfitting, ultimately leading to improved performance under unseen conditions.

### v3

@@ Abstract
Concrete workability is essential for construction quality, with the slump test being the most widely used on-site method for its assessment. However, traditional slump testing is manual, time-consuming, and highly operator-dependent, making it unsuitable for continuous or real-time monitoring during placement. To address these limitations, we present SlumpGuard, an AI-powered vision system that analyzes the natural discharge flow from a mixer-truck chute using a single fixed camera. The system performs automatic chute detection, pouring-event identification, and video-based slump classification, enabling quality monitoring without sensors, hardware installation, or manual intervention. We introduce the system design, construct a site-replicated dataset of over 6,000 video clips, and report extensive evaluations demonstrating reliable chute localization, accurate pouring detection, and robust slump prediction under diverse field conditions. An expert study further reveals significant disagreement in human visual estimates, highlighting the need for automated assessment. Demonstration videos are available at this URL.
@@ 1 Introduction
Concrete stands as the backbone of modern construction infrastructure, with over 14 billion cubic meters produced annually worldwide, making it the most widely consumed material after water [hoang2016:estimating]. Among the critical properties governing concrete performance, workability, which is primarily assessed through slump testing, represents a fundamental parameter that directly influences mixing efficiency, transportation logistics, placement operations, and compaction effectiveness on construction sites [chen2022:research, ojala2024:estimating]. However, the complex and dynamic nature of construction environments presents unprecedented challenges for maintaining consistent concrete quality control. In particular, conventional manual slump testing, still the industry standard, often fails to meet the demanding requirements of large-scale, time-sensitive projects.
The construction industry therefore continues to struggle with real-time quality assurance. On-site inspection of fresh concrete is typically performed only intermittently rather than for every delivered batch. For example, the Korean Construction Specification KCS 14 20 10 mandates slump inspection only once per day or once every 120 cubic meters (approximately 20 truckloads), making continuous, batch-level verification infeasible in practice. As a result, quality deviations that arise during mixing, transportation, or placement may remain undetected until concrete is already being poured, leading to delays, increases in cost, and inconsistency due to human measurement variability [tuan2021:situ].
To address these challenges, researchers have increasingly explored automated and data-driven methods for assessing fresh concrete properties. As part of these efforts, sensor-based approaches employ piezoelectric, depth, and fiber-optic measurements to quantify workability-related characteristics. Leveraging these sensing modalities, recent work has explored combining the captured signals with deep learning and IoT-based frameworks to enable automated analysis of fresh concrete properties [han2025:field, kim2018:visualization, govindaraju2025:real]. However, sensor-based solutions heavily rely on hardware installation and calibration, and their practicality is often constrained by on-site regulatory restrictions and variations in equipment conditions across different construction environments.
To overcome these limitations, visual information–based methods for slump estimation have also been proposed. Yucelai et al. [yucelai:ai] introduced a stereo-vision system for predicting slump during laboratory slump tests. Other studies attempted to estimate slump by installing cameras in proximity to the concrete mixer and analyzing captured images [idrees2024:automatic, hao2025:concrete]. Despite their promise, these approaches face significant barriers to real-world deployment. The method of Yucelai et al. [yucelai:ai] requires repeated on-site slump testing, which is rarely performed during routine concrete placement, limiting its practical applicability. Methods based on mixer-mounted cameras [idrees2024:automatic, hao2025:concrete] are also susceptible to dust and slurry contamination, leading to frequent maintenance and reduced operational reliability. Moreover, image-only frameworks generally fail to capture the flow-driven characteristics of fresh concrete, making it difficult to reflect the inherent dynamic behavior underlying slump.
To address the limitations of previous automated approaches, we introduce SlumpGuard, a fully autonomous, camera-only vision system designed for real-site deployment. Operating with a single fixed camera that analyzes the natural discharge flow, SlumpGuard avoids hardware installation and contamination issues and integrates seamlessly into existing workflows. This lightweight, non-intrusive design enables continuous, batch-level monitoring under diverse on-site conditions. At its core, SlumpGuard is built upon a modular three-stage pipeline that structures the overall task into interpretable components.
To operationalize this goal in real construction environments, SlumpGuard decomposes the problem into three stages, each addressing a distinct challenge encountered during concrete placement. First, the system performs robust chute detection to automatically localize the discharge chute in unconstrained on-site videos. This step isolates the region where concrete flows, enabling stable analysis regardless of camera placement, background clutter, or chute rotation. Second, SlumpGuard identifies both when and from which chute concrete begins to fall by analyzing pixel-level motion using optical flow. This automated detection of flow onset ensures that the system processes only frames containing concrete discharge, eliminating dependence on manual trimming or operator input. Third, once the relevant video segment is identified, a video-based slump classification module analyzes the temporal evolution of concrete flow to estimate the slump category in real time. By repeatedly predicting slump over successive frame windows and aggregating the results, the system achieves stable and reliable batch-level slump assessment. This modular design enhances interpretability and facilitates adaptation to different site conditions, while the motion-gated processing ensures real-time operation by restricting inference to concrete–flow intervals.
To assess the practical effectiveness of this three-stage pipeline and the deployability of SlumpGuard, we conducted extensive experiments on more than 6,000 video clips collected under site-replicated conditions. Our evaluation shows that each module contributes complementary strengths toward stable real-world operation. The chute detection module (Stage 1) consistently identified mixer-truck chutes across diverse geometric and illumination conditions, while also achieving high reliability in recognizing which chute was actively discharging concrete (Stage 2). The video-based slump classification module (Stage 3) further demonstrated strong performance, achieving over 80% accuracy despite substantial visual variability in concrete flow. This variability underscores the practical need for objective, automated systems capable of consistent assessment in conditions where human judgment is inherently limited. Together, these findings indicate that SlumpGuard offers a reliable and field-ready vision-based alternative to conventional manual or sensor-dependent inspection methods.
@@
The remainder of this paper is organized as follows. Section 2 reviews existing studies on concrete workability assessment and automated monitoring approaches. Section 3 describes the dataset collection process and annotation methodology. Section 4 details the architecture of the proposed SlumpGuard system, including its three-stage pipeline and data augmentation strategies. Section 5 presents experimental results and performance evaluation under real-site conditions. Section 6 introduces how SlumpGuard is deployed and utilized in real construction sites. Section 7 and Section 8 provide a discussion of SlumpGuard, including its implications and limitations. Finally, Section 9 concludes the paper and outlines future research directions.
@@ 2.1 Optical Flow
Optical flow is a computer vision technique used to estimate the motion of objects or scenes across video frames by analyzing the movement of pixels. It represents motion as a vector field, where each vector indicates the displacement of a pixel between consecutive frames. There are two main approaches: dense optical flow, which computes motion for every pixel, and sparse optical flow, which estimates motion only at selected feature points. While dense optical flow [horn1981:determining] offers a detailed view of the entire scene’s motion, it is computationally intensive. Sparse optical flow is more efficient but may overlook fine-grained motion.
In this work, we adopt the Lucas-Kanade method [lucas1981:iterative], a widely used algorithm for sparse optical flow, suitable for real-time applications such as tracking concrete flow. It assumes that the intensity of a point remains constant over time:
@@
FlowNet [dosovitskiy2015:flownet] pioneered the use of convolutional neural networks for end-to-end optical flow prediction, learning motion patterns directly from synthetic datasets. Building on this idea, FlowNet2 [ilg2017:flownet], PWC-Net [sun2018:pwc], and LiteFlowNet [hui2018:liteflownet] introduced architectural improvements such as pyramid-based warping, cost volume reasoning, and hierarchical refinement. More recently, RAFT [teed2020:raft] achieved significant performance gains by leveraging dense, all-pairs correlations and iterative updates to produce highly accurate flow estimates.
Previous studies have applied optical flow combined with imaging techniques for slump prediction under controlled experimental settings [tuan2021:situ, kim2018:visualization]. In contrast, our approach is specifically designed for robust, real-world deployment on construction sites, enabling continuous, automated slump quality monitoring under diverse and dynamic field conditions.
@@ 2.2 Object Detection
Object detection is a fundamental computer vision task widely applied in construction to automatically identify and localize key entities such as workers, equipment, and materials in images or videos. Recent studies have demonstrated its effectiveness in construction site monitoring [pfitzner2025:monitoring, kim2023:small].
State-of-the-art detection frameworks such as Faster R-CNN [ren2015:faster], Mask R-CNN [he2017:mask], SSD [liu2016:ssd], and YOLO [redmon2016:you] offer a good balance between speed and accuracy, making them well-suited for real-time applications on construction sites.
However, in construction environments where objects like materials and equipment often appear at arbitrary angles, standard axis-aligned bounding boxes may be insufficient. In such cases, oriented object detection becomes essential. Early approaches like R2CNN [jiang2017:r2cnn] introduced rotation-invariant proposals, while more recent models such as Oriented R-CNN [xie2021:oriented] and ReDet [han2021:redet] employ rotation-aware architectures to improve detection accuracy. These methods are particularly effective for automated inspection tasks involving rotated or non-axis-aligned objects.
@@ 2.3 Video Classification
Video classification is a fundamental task in computer vision that involves assigning semantic labels to video clips by analyzing both spatial and temporal information inherent in the data. Recent advances in this domain have been largely driven by two primary classes of deep learning architectures: 3D convolutional neural networks (3D CNNs) and vision transformers [dosovitskiy2020:image] tailored for video.
3D CNNs like ResNet3D [tran2018:closer] extend 2D convolutions to the temporal dimension, effectively capturing spatiotemporal features and achieving strong results on benchmarks such as Kinetics-400 [kay2017:kinetics]. Transformer-based models, notably TimeSformer [bertasius2021:space], use divided spatial and temporal self-attention to model long-range dependencies efficiently, surpassing many convolutional approaches. Further advancements include VideoSwin [liu2022:video] and MViT [fan2021:multiscale], which incorporate hierarchical and multiscale attention mechanisms to improve video representation learning and classification performance.
These video classification models have also been increasingly applied in the construction industry, playing a pivotal role in streamlining quality supervision, activity recognition, and concrete workability assessment. Previous studies have also explored video-based approaches for material quality monitoring and evaluation [guo2025:egocentric, ojala2024:estimating, idrees2024:automatic], alongside behavior classification systems aimed at improving worker safety and productivity [li2022:action, yang2023:transformer].
@@ 3.1 Data Collection
In the experimental process, considering cost efficiency, we strategically utilized a single stereo camera for video acquisition. The stereo camera allowed simultaneous capturing of video footage from two distinct viewpoints in a single filming session, allowing richer data collection within a limited number of experiments. All videos were recorded at 30 frames per second (FPS), with a stereo capture resolution of $3840\times 1080$ composed of two individual camera views, each operating at $1920\times 1080$. This approach maximized data efficiency while clearly capturing the concrete flow through the chute. Additionally, we performed traditional slump cone measurements at regular intervals, establishing precise ground-truth labels corresponding to each video frame. This example is shown in Figure 1.
@@ 3.2 Preprocessing and Annotation
Since we employed a stereo camera setup, the collected videos consisted of two separate streams. Therefore, as part of preprocessing, the original stereo data was divided into two individual video sequences, effectively doubling the amount of data for analysis. Considering that concrete discharges directly from the ready-mixed concrete chute, our preprocessing specifically targeted the chute region. Bounding box annotations were performed by categorizing the chute region into two classes, as shown in Figure 2: an “Unrotated Chute” and a “Chute”. The “Unrotated Chute” corresponds to an axis-aligned bounding box that fully encloses the chute, ensuring that both endpoints are contained within the box. In contrast, the “Chute” class uses rotated bounding boxes to precisely capture the chute’s orientation.
@@ 4.1 Chute Detection
To automatically identify when concrete is falling from the chute, we first detect the chute region in the video using an object detection model (see Step 1 in Figure 3). We employ YOLOv8 [yolov8], a state-of-the-art object detection model, known for its fast inference performance and high accuracy, making it suitable for real-time object detection. We initialize the model with weights pretrained on the DOTA dataset [xia2018:dota], which is designed for detecting oriented objects in aerial imagery and is well-suited for our task of detecting rotated chutes. Using this model, we simultaneously predict both axis-aligned and rotated-bounding boxes for the chutes, enabling us to detect the chute area in pouring videos. The detected chute is represented by a bounding box with parameters $(x,y,w,h,\theta)$, indicating the center coordinates, width, height, and rotation angle of the box.
@@ 4.2 Detecting Concrete Placement Location and Timing
On construction sites, concrete is discharged from two separate chutes connected to individual concrete mixer trucks. In this study, we detect which chute the concrete is falling from and accurately identify the starting point of the flow. This step is essential to ensure that the subsequent video classification model receives only the segments where concrete is actually flowing, enabling more precise slump analysis (see Step 2 in Figure 3). Since video segments without concrete flow are not relevant for analysis, filtering out such irrelevant data in advance is a critical preprocessing step.
To precisely determine when and from which chute the concrete starts to fall, we perform motion analysis based on the segmented chute videos $V^{L}$ and $V^{R}$ obtained in Section 4.1. For each video segment, we extract bounding boxes for the chutes from the object detector and compute optical flow at the center point of each bounding box. Specifically, we apply the Lucas-Kanade method, a sparse optical flow algorithm, to track the motion vector $(u,v)$ at the center $(x,y)$ across consecutive frames. When the tracked center point crosses the bottom boundary of the bounding box over time, we identify the moment of concrete drop as well as the specific chute through which the concrete is discharged.
@@ 4.2.2 Determining the Location and Timing
To make the temporal change explicit, we define the center point of the tracked chute in each frame as $(x_{t},y_{t})$ at time $t$, and as $(x_{t-1},y_{t-1})$ at the previous frame t1t-1. The optical flow vector $(u,v)$ represents the displacement from frame t1t-1 to $t$, so that $(x_{t},y_{t})=(x_{t-1}+u,y_{t-1}+v)$. The endpoints of the bottom edge of the rotated bounding box, $\mathbf{p_{1}}=(x_{1}+\alpha,y_{1}+\alpha)$ and $\mathbf{p_{2}}=(x_{2}+\alpha,y_{2}+\alpha)$, are computed. This translation mitigates false triggering caused by optical-flow sensitivity to shadows by slightly raising the decision boundary for a more stable crossing detection. We empirically use $\alpha=100$, applied as $\alpha$ for the right chute and $\alpha$ for the left chute.
@@ 4.3 Slump Prediction and Anomaly Detection
To enable efficient and real-time estimation of concrete slump, we leverage the drop timing and location identified in Section 4.2. Specifically, given the drop time $t$, we predict the slump using video data from the corresponding chute region. This process corresponds to Step 3 of the overall pipeline shown in Figure 3. To focus on the continuous flow of concrete, we adopt a video classification model, ResNet 3D [tran2018:closer]. Let $V$ denote the chute video (i.e., $V^{R}$ or $V^{L}$), and $t$ the detected drop time.
@@ 4.3.2 Data Augmentation
Given that construction sites are exposed to a wide range of lighting and environmental conditions, data augmentation is essential to improve the model’s generalization performance. To simulate these real-world variations, we employ a set of augmentation techniques, including ColorJitter, gamma correction, horizontal flipping, and contrast adjustment. We also apply MixUp [zhang2017:mixup] and CutMix [yun2019:cutmix], which blend either the pixel content or the labels of two training examples to generate mixed samples. While MixUp interpolates both images and labels, CutMix replaces a region of an image with a patch from another image, adjusting the labels proportionally to the area. These strategies encourage the model to learn more robust and smoother decision boundaries, and to be less sensitive to spurious correlations between local image regions and class labels.
Together, these augmentations help the model handle diverse visual appearances, such as changes in brightness, shadows, occlusion, and orientation, that are commonly encountered in construction site environments. They also promote better regularization by expanding the training distribution and reducing overfitting, ultimately leading to improved performance under unseen conditions.